OpenStudy (anonymous):

determining the domain of a function. s=1/3t+12

OpenStudy (campbell_st):

is the equation \[s = (1/3) t + 12\] or \[s = 1/(3t) + 12\]

OpenStudy (rulnick):

Or even 1/(3t+12)??

OpenStudy (campbell_st):

very true

OpenStudy (anonymous):

no it is the first one Itook it straight from the book

OpenStudy (campbell_st):

so the domain is all real t all real t except t = 0 or all real t except t = -4

OpenStudy (rulnick):

In the first case it's "all real numbers," in the second it's "all real numbers except 0," and in the third it's "all real numbers except -4."

OpenStudy (campbell_st):

well its all real t...

OpenStudy (rulnick):

Yup, the first case then.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do i write that in interval notation?

OpenStudy (campbell_st):

\[(-\infty, \infty)\]

OpenStudy (anonymous):

\[s(t)=t/3+12\] Domain: Set of all real numbers. Interval notation: \[(-\infty,\infty)\]